Anthony Tedesco sold his used mountain bike and accessories for $308. If he received  six times as much money for the bike as he did for the accessories, find how much money he received for the bike.
Answer by ikleyn(52778)

.

    x + 6x = 308

      7x   = 308

       x   = 308/7 = 44.


ANSWER.  44*6 = 264 dollars for the bike.

Solved.